Casey Nelson is a divisional manager for Pigeon Company. His annual pay raises are largely determined by his divisions returReq 1 Reg 2 Req3 Req 4A Req 4B What is the projects internal rate of return? (Round your answer to whole decimal place i.e.Req1 Req2 Req3 Req 4A Req 4B What is the projects simple rate of return? (Round percentage answer to 1 decimal place.) SimplWould the company want Casey to pursue this investment opportunity? Yes IONO < Rea 3 < Reg 3 Req4B > Reg ABWould Casey be inclined to pursue this investment opportunity? Yes Coro No ( Req 4A Reg 4B)
Answer #1
Net operating income 430000
Add: Depreciation 720000
Net cash flows 1150000
1
Now Year 1 Year 2 Year 3 Year 4 Year 5
Investment cost -3600000
Net cash flows 1150000 1150000 1150000 1150000 1150000
Total cash flows -3600000 1150000 1150000 1150000 1150000 1150000
PV factor @ 16% 1 0.862 0.743 0.641 0.552 0.476
Present value of cash flows -3600000 991300 854450 737150 634800 547400
Net present value 165100
2
PV factor internal rate of return=3600000/1150000 = 3.130
The PV factor 3.130 for 5 years is closest to 18%
Internal rate of return = 18%
3
Simple rate of return = Net operating income/Investment cost
Simple rate of return = 430000/3600000= 11.9%
4a
Yes, the company would want Casey to pursue this investment as Net Present value is positive
4b
No, Casey would not be inclined to pursue this investment as as his ROI will decrease